OpenStudy (anonymous):

If 2x = 32, find the value of 10 - 4*x

OpenStudy (anonymous):

Well, 2x=32 so x=32/2 x=16 Now can you find the answer?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhh i see where i messed up i used times

OpenStudy (texaschic101):

2x = 32 x = 32/2 x = 16 10 - 4x 10 - 4(16) 10 - 64 = - 54 answer is -54

OpenStudy (anonymous):

Good @texaschic101!

OpenStudy (texaschic101):

thank you Chineseboy :)

OpenStudy (anonymous):

yes thank u i have another one its so confusing

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

What it is?

OpenStudy (anonymous):

Thank you @texaschic101!

OpenStudy (texaschic101):

no problem....you gave the correct way to solve it....I gave the answer...was not supposed to do that